OpenStudy (anonymous):

If a/4 = 9/a, then A)a²=36 B)4a = 9a C)a + 4 = 9 + a D)a - 4 = 9 - a

OpenStudy (aravindg):

whats so difficult here?

Parth (parthkohli):

Let's see what you get once you multiply both sides through \(4a\).

OpenStudy (aravindg):

also if a/b=c/d ad=bc try to use that

OpenStudy (anonymous):

So it's B?

Parth (parthkohli):

@themonkey Show your work please?

OpenStudy (anonymous):

4a^2/16a 36a/4a^2